One canned juice drink is 15% lemon juice; another is 5% lemon juice. How many liters of each should be mixed together in order to get 10L that is 10% lemon juice?
Answer by jorel1380(3719)

10Lx.10=1 Liter pure juice
p+q=10
.15p+.05q=1
15p+5q=100
5p+5q=50
10p=50
p=5
q=5
You need 5 liters of each for the mix..